OpenStudy (anonymous):

Suppose that the functions f and g are defined for all real numbers x as follows. f(x)=x^3 g(x)=4x^2. Write the expressions for (f*g)(x) and(f+g)(x) and evaluate (f-g)(-1)

OpenStudy (anonymous):

@tkhunny @dan815

OpenStudy (tkhunny):

These are just notation exercises. There is only learning what the notation means. (f*g)(x) = f(x) * g(x) (f+g)(x) = f(x) + g(x) (f-g)(-1) = f(x) - g(x) Build those and substitute any values shown.